If Ram walks from his house at 5kmph, he reaches school 5 min early & if he walks at 4 kmph he reaches 5 min late. What is the actual distance b/w Ram's house & his school?
Answer by rothauserc(4718)

we use rate * time(t) = distance
there are two equations which we set equal to each other
5*(t-5) = 4*(t+5)
5t -25 = 4t +20
t = 45 minutes
distance between Ram's home and his school
distance = 5 * ((45-5)/60) = 3.333333333 km